OpenStudy (anonymous):

PLSE HELPPPPPPPPPPP A solid sphere of diameter 8 cm is melted and drawn into a cylindrical wire of uniform width.If the length of the wire is 12 m, find its width.

OpenStudy (shubhamsrg):

Equate volumes.

Directrix (directrix):

The volume of the sphere is given by V = (4/3)*pi*r^3 where r = 4. V = (4/3)*pi*(4)^3 --> Volume of Sphere

Directrix (directrix):

The volume of the cylindrical wire of uniform width is given by V = pi*r^2 * h where h = 12. V = pi*r^2 * (12) --> Volume of Cylinder

Directrix (directrix):

Volume of Sphere = Volume of Cylinder (4/3)*pi*(4)^3 = pi*r^2 * (12) Solve this equation for r, the radius of the base of the cylinder. Double r to get the width of the wire. @sumitra